# Improving Outdoor Multi-cell Fingerprinting-based Positioning via Mobile Data Augmentation

### v1

@@ II-C Radio Features Augmentation via Non-Parametric Regression
where each $\hat{\mathbf{r}}_{j}$ is a measurement vector recorded at the known location $\hat{\mathbf{p}}_{j}$. Given a query spatial location $\mathbf{p}\in\mathbb{R}^{2}$ (either sampled from a test set, sampled from the KDE in Sec. II-B, or otherwise selected), our goal is to synthesize a realistic multi-cell radio fingerprint ${\mathbf{r}}(\mathbf{p})=[r_{1}(\mathbf{p}),\dots,r_{N_{c}}(\mathbf{p})]^{\top}$ consisting of per-PCI RSRP measurements that are consistent with the empirical database $\hat{\mathbf{R}}$.
@@ II-D Analysis and Benefits
    The pipeline is essentially training-free for radio-feature synthesis. KDE sampling and the nearest-neighbor algorithm avoid iterative, parameter-intensive training loops typical of RF, GPR or deep models. The practical consequences are significant: lower computational and engineering costs, faster time-to-deployment, and elimination of long retraining cycles whenever new MDT records become available. Because no complex model parameters must be learned and propagated, newly collected data can be appended and used immediately to refresh the empirical database or to re-sample the KDE, providing a seamless data-integration path that directly benefits both augmentation and downstream positioning.
@@
    The design supports privacy- and data-minimization strategies. Since the spatial model can be represented by a compact density parameterization (or by synthetic locations sampled from it), operators can avoid sharing raw, user-level MDT traces across different sites. Moreover, in distributed deployments the detailed per-PCI measurements can remain in the data domain where they were collected, reducing privacy exposure and simplifying compliance with data-protection requirements.
@@
III Fingerprinting-based Positioning and
@@
III-A Multi-cell Fingerprinting-based Positioning
There exist several neural network-based approaches to multi-cell fingerprinting-based positioning, including those in [34, 35]. Other works, including [2, 5], rely on non-parametric predictors such as weighted k‐nearest‐neighbor (wKNN) for determining position information. In this work, we leverage wKNN as an illustrative example of a positioning algorithm that can possibly benefit from mobile data augmentation.
In multi-cell fingerprinting-based positioning, a set of base stations (cells) act as anchors whose received‐signal measurements are used to infer an unknown user location. As introduced in Sec. II-C, we denote by $\hat{\mathbf{R}}$ the reference fingerprint database, comprising $m$ radio measurements with respect to $N_{c}$ anchors and user position information. For the ii‑th user (at unknown position $\mathbf{p}_{i}\in\mathbb{R}^{2}$), we observe a radio‑measurement vector
@@ III-A Multi-cell Fingerprinting-based Positioning
and we perform neighbor selection computing all distances $d\bigl(\mathbf{r}_{i},\hat{\mathbf{r}}_{j}\bigr)$ for all $j=1,\dots,m$. We denote the indices of the $K$ reference fingerprints with smallest distance to $\mathbf{r}_{i}$ as
@@
III-B MDT-based Reference Scenarios
In this work, we leverage a real-world MDT dataset collected across the city of Bologna (Italy) and provided by an Italian MNO. In the dataset, each sample corresponds to a user record, including its latitude, longitude, and RSRP measurements that are reported for all visible PCIs within its reception range. The measurements cover four distinct regions, namely a dense urban area, mixed urban-rural zones, an airport, and a highway. To study the impact of augmentation under diverse environmental and operational conditions, the four scenarios are chosen to be heterogeneous.
These regions vary in spatial density, mobility profiles, and radio coverage characteristics, offering a representative basis for evaluating RSRP-based fingerprinting positioning systems. To reflect realistic propagation environments, each area is modeled under predominantly non-line-of-sight (NLOS) conditions, which are common in urban and suburban deployments. Additionally, a shadow fading term is introduced for each region following the 3GPP TR 38.901 V18 [33] standard and incorporated into the PCI-wise radio feature modeling process. Visual representations of these regions are provided in Fig. 2. Their quantitative characteristics, including spatial extent, number of UEs, user density, and shadow fading variance $(\sigma_{s}^{2})$, are summarized in Tab. II.
@@ III-B MDT-based Reference Scenarios
The selected areas include:
  • The city center area, which represents the most compact and dense region and covers the historical core of Bologna and served by 140 distinct PCIs.
IV Numerical Results
This section evaluates the proposed mobile data augmentation architecture both stage-by-stage and end-to-end, emphasizing its effect on fingerprinting-based positioning. We first validate each module in isolation to verify that the Spatial Features Augmentation and the Radio Features Augmentation behave as intended (Sec. IV-A and Sec. IV-B, respectively). Then, we assess the impact of the full augmentation pipeline on multi-cell fingerprinting-based positioning across the four real-world MDT scenarios introduced in Sec. III-B (Sec. IV-C).
Our primary objective is to quantify improvements in positioning performance, while also ensuring that intermediate modules produce statistically faithful and useful synthetic data. Accordingly, module evaluations report relevant per-module metrics (i.e., multivariate two-sample KS tests and radio prediction error metrics such as MAE), whereas the end-to-end study focuses on positioning accuracy. All results compare against established baselines for the respective tasks, while Sec. IV-D investigates the scaling effect of progressively augmenting the reference database (i.e., how adding more synthetic samples affects positioning).
IV-A Results on Spatial Features Augmentation
@@
To assess whether synthetic samples are statistically similar to real MDT locations we employ a multivariate two-sample KS test. Using this test, we assess whether synthetic user positions generated by KDE and the competing methods are statistically consistent with the spatial distribution of positions in the original test set. Under the null hypothesis the real samples and synthetic samples originate from the same distribution; we reject the null whenever the KS $p$-value falls below the significance level $\alpha=0.05$. Tab. III reports the average test statistic and the average $p$-value for each model and region over ten independent runs, with associated 95% confidence interval (CI). Fig. 3 provides an illustrative example of real and synthetic samples for one representative region.
@@
IV-B Results on Radio Features Augmentation
Analogously, for radio feature augmentation we compare four methods for synthesizing per-PCI RSRP fingerprints [19]: (i) the proposed spatial KNN-based augmentation with additive log-normal shadowing; (ii) a RF regression trained per-PCI; (iii) a GPR with a SE kernel; and (iv) a GPR with a RQ kernel. All radio feature models are trained on region-specific training sets and evaluated on held-out test sets using the MAE on per-PCI RSRP values.
This evaluation is conducted independently for each geographic region. The reported values in Tab. IV represent the average MAE over ten independent runs, with associated 95% CI.
The results demonstrate that all models achieve relatively low MAE values, indicating effective RSRP estimation across diverse urban and semi-urban regions. The KNN model consistently yields the lowest MAE across all regions, particularly excelling in the airport and city center zones, which contain the highest number of UEs. In contrast, the GPR models show slightly worse performance overall. The RQ variant consistently outperforms the standard GPR across all regions, as it incorporates a more flexible kernel capable of modeling multi-scale signal variations. RF exhibits moderate performance, with accuracy generally falling between KNN and GPR.
@@ IV-B Results on Radio Features Augmentation
Table IV: MAE of RSRP prediction for KNN, RF, and GPR with SE and RQ kernels in each reference region.
IV-C Results on Multi-cell Fingerprinting-based Positioning
In this section, we evaluate the end-to-end impact of the proposed mobile data augmentation architecture on multi-cell fingerprinting-based positioning. Starting from the region-specific reference database, we compare positioning performance obtained with the original operator-collected fingerprints against databases augmented with synthetic samples produced by the spatial and radio augmentation stages. The positioning algorithm under test is the wKNN estimator introduced in Sec. III-A; results are reported for all four MDT reference scenarios.
We compare our proposed two-stage architecture KDE-KNN against three alternative two-stage configurations: KDE-RF and KDE-GPR, using both SE and RQ kernels. For each MDT scenario, we evaluate a set of augmentation rates $A\in\{1,5,10,20,30\}$, where $A$ denotes the multiplicative increase of the fingerprinting database relative to the original ($A=1$ corresponds to no augmentation). For every combination of scenario, augmentation rate and architecture, we perform ten independent runs. In the remanding of this section, we report average MAE values for positioning together with the 95% CI.
@@ IV-C Results on Multi-cell Fingerprinting-based Positioning
    City Center Area (Tab. VI). In the city center scenario, which features the highest measurement density and the smallest spatial extent, the wKNN estimator applied to the original MDT baseline attains a mean positioning error of 20.0820.08\,m. With the proposed KDE-KNN pipeline and an augmentation factor $A=20$, the mean error decreases to 17.8017.80\,m, corresponding to an $\approx 11.4\%\$relative improvement. Although the absolute reduction ($\approx 2.3\,$m) is modest, it is noteworthy given the strong baseline performance. Moreover, KDE-KNN consistently outperforms the alternative two-stage configurations, which yield smaller or no gains; in particular, the KDE-RF variant may be detrimental, since RF-based RSRP predictions can produce augmented fingerprints that degrade positioning accuracy.
@@
    Stadium Area (Tab. V).
    Augm. rate $A$ KDE-KNN KDE-RF KDE-GPR KDE-GPR (RQ)
@@
    Table V: Positioning performance of KDE-KNN and baselines across different augmentation rates in stadium scenario.
    In the stadium scenario, characterized by a high measurement density but a very confined spatial extent and an intermediate number of PCIs, the proposed KDE-KNN pipeline yields substantial gains. With augmentation factor $A=20$, the mean positioning error drops from $36.70$ m to $28.32$ m, an absolute improvement of $8.38$ m and a relative reduction of $\approx 23\%$. By contrast, the alternative two-stage configurations provide only marginal or no improvements across augmentation levels.
  3. 3.
    Airport Area (Tab. VII).
    Augm. rate $A$ KDE-KNN KDE-RF KDE-GPR KDE-GPR (RQ)
@@
    Table VI: Positioning performance of KDE-KNN and baseline models across different augmentation rates in the city center scenario.
    Augm. rate $A$ KDE-KNN KDE-RF KDE-GPR KDE-GPR (RQ)
@@
    Table VII: Positioning performance of KDE-KNN and baseline models across different augmentation rates in the airport scenario.
@@
    Augm. rate $A$ KDE-KNN KDE-RF KDE-GPR KDE-GPR (RQ)
@@
    Table VIII: Positioning performance of KDE-KNN and baseline models across different augmentation rates in the highway scenario.
    In the highway scenario, where UE measurements are highly concentrated along the linear roadway (i.e., samples cluster on the highway) but sparse elsewhere across the largest geographic extent and subject to high mobility, augmentation delivers the largest benefit. With the proposed KDE-KNN pipeline and augmentation factor $A=30$, the mean positioning error falls from $115.51$ m to $79.82$ m, a relative reduction of $\approx 30\%$. This substantial improvement indicates that realistic augmentation can compensate for the anisotropic and sparse spatial sampling typical of highway measurements; by contrast, the alternative two-stage configurations produce markedly smaller gains and are consistently outperformed by KDE-KNN across augmentation levels.
These findings lead to three relevant conclusions: (i) the KDE-KNN pipeline consistently improves positioning accuracy across all evaluated scenarios, with the largest relative gains in spatially sparse or structurally complex regions; (ii) the alternative two-stage baselines fail to match the proposed approach across the tested augmentation rates; and (iii) performance gains saturate beyond a region-specific augmentation factor, so adding more synthetic samples does not by itself guarantee better accuracy, and poorly matched synthetic fingerprints can even degrade it.
IV-D Statistical Model Comparison
and their sample variance as

### v2

@@ IV-C Radio Features Augmentation via Non-Parametric Regression
where each $\hat{\mathbf{r}}_{j}$ is a measurement vector recorded at the known location $\hat{\mathbf{p}}_{j}$. Given a query spatial location $\mathbf{p}\in\mathbb{R}^{2}$ (either sampled from a test set, sampled from the KDE in Sec. IV-B, or otherwise selected), our goal is to synthesize a realistic multi-cell radio fingerprint ${\mathbf{r}}(\mathbf{p})=[r_{1}(\mathbf{p}),\dots,r_{N_{c}}(\mathbf{p})]^{\top}$ consisting of per-PCI RSRP measurements that are consistent with the empirical database $\hat{\mathbf{R}}$.
@@ IV-D Analysis and Benefits
    The pipeline is essentially training-free for radio-feature synthesis. KDE sampling and the nearest-neighbor algorithm avoid iterative, parameter-intensive training loops typical of RF, GPR, or deep models. The practical consequences are significant: lower computational and engineering costs, faster time-to-deployment, and elimination of long retraining cycles whenever new MDT records become available. Because no complex model parameters must be learned and propagated, newly collected data can be appended and used immediately to refresh the empirical database or to re-sample the KDE, providing a seamless data-integration path that directly benefits both augmentation and downstream positioning.
@@
    The design supports privacy- and data-minimization strategies. Since the spatial model can be represented by a compact density parameterization (or by synthetic locations sampled from it), operators can avoid sharing raw, user-level MDT traces across different sites. Moreover, in distributed deployments, the detailed per-PCI measurements can remain in the data domain where they were collected, reducing privacy exposure and simplifying compliance with data-protection requirements.
@@
V Fingerprinting-based Positioning and
@@
V-A Multi-cell Fingerprinting-based Positioning
There exist several neural network-based approaches to multi-cell fingerprinting-based positioning, including those in [14, 22]. Other works, including [16, 7, 8], rely on non-parametric predictors such as weighted k‐nearest‐neighbor (wKNN) for determining position information. In this work, we leverage wKNN as an illustrative example of a positioning algorithm that can possibly benefit from mobile data augmentation.
In multi-cell fingerprinting-based positioning, a set of base stations (cells) act as anchors whose received‐signal measurements are used to infer an unknown user location. As introduced in Sec. IV-C, we denote by $\hat{\mathbf{R}}$ the reference fingerprint database, comprising $m$ radio measurements with respect to $N_{c}$ anchors and user position information. For the ii‑th user (at unknown position $\mathbf{p}_{i}\in\mathbb{R}^{2}$), we observe a radio‑measurement vector
@@ V-A Multi-cell Fingerprinting-based Positioning
and we perform neighbor selection computing all distances $d\bigl(\mathbf{r}_{i},\hat{\mathbf{r}}_{j}\bigr)$ for all $j=1,\dots,m$. We denote the indices of the $K$ reference fingerprints with the smallest distance to $\mathbf{r}_{i}$ as
@@
V-B MDT-based Reference Scenarios
In this work, we leverage a real-world MDT dataset collected across the city of Bologna (Italy) and provided by an Italian MNO. In the dataset, each sample corresponds to a user record, including its latitude, longitude, and RSRP measurements that are reported for all visible PCIs within its reception range. To ensure the reliability of these spatial labels subject to GPS multipath and drift, we applied a quality control filter based on the positioning uncertainty fields defined in 3GPP standards. Specifically, we utilized the uncertainty semi-major axis to discard samples with a positioning uncertainty radius exceeding 10 m. This pre-processing step effectively mitigates the impact of location drift, ensuring that the ground-truth coordinates used for training remain robust even in challenging scenarios. The measurements cover four distinct regions, namely a dense urban area, mixed urban-rural zones, an airport, and a highway. To study the impact of augmentation under diverse environmental and operational conditions, the four scenarios are chosen to be heterogeneous.
These regions vary in spatial density, mobility profiles, and radio coverage characteristics, offering a representative basis for evaluating RSRP-based fingerprinting positioning systems. To reflect realistic propagation environments, each area is modeled under predominantly non-line-of-sight (NLOS) conditions, which are common in urban and suburban deployments. Additionally, a shadow fading term is introduced for each region following the 3GPP Release 10 [4] standard and incorporated into the PCI-wise radio feature modeling process. Visual representations of these regions are provided in Fig. 2. Their quantitative characteristics, including spatial extent, number of UEs, user density, and shadow fading variance $(\sigma_{s}^{2})$, are summarized in Tab. II.
@@ V-B MDT-based Reference Scenarios
The selected areas include:
  • The city center area, which represents the most compact and dense region and covers the historical core of Bologna and is served by 140 distinct PCIs.
VI Numerical Results
This section evaluates the proposed mobile data augmentation architecture both stage-by-stage and end-to-end, emphasizing its effect on fingerprinting-based positioning. We first validate each module in isolation to verify that the Spatial Features Augmentation and the Radio Features Augmentation behave as intended (Sec. VI-A and Sec. VI-B, respectively). Then, we assess the impact of the full augmentation pipeline on multi-cell fingerprinting-based positioning across the four real-world MDT scenarios introduced in Sec. V-B (Sec. VI-C).
Our primary objective is to quantify improvements in positioning performance, while also ensuring that intermediate modules produce statistically faithful and useful synthetic data. Accordingly, module evaluations report relevant per-module metrics (i.e., multivariate two-sample KS tests and radio prediction error metrics such as MedAE), whereas the end-to-end study focuses on positioning accuracy. All results compare against established baselines for the respective tasks, while Sec. VI-D investigates the scaling effect of progressively augmenting the reference database (i.e., how adding more synthetic samples affects positioning).
VI-A Results on Spatial Features Augmentation
@@
To assess whether synthetic samples are statistically similar to real MDT locations we employ a multivariate two-sample KS test. Using this test, we assess whether synthetic user positions generated by KDE and the competing methods are statistically consistent with the spatial distribution of positions in the original test set. Under the null hypothesis, the real samples and synthetic samples originate from the same distribution; we reject the null whenever the KS $p$-value falls below the significance level $\alpha=0.05$. Tab. III reports the average test statistic and the average $p$-value for each model and region over ten independent runs, with associated 95% confidence interval (CI). Fig. 3 provides an illustrative example of real and synthetic samples for one representative region.
@@
VI-B Results on Radio Features Augmentation
Similarly, for radio feature augmentation we compare eight methods for synthesizing per-PCI RSRP fingerprints [39]: (i) the proposed spatial KNN-based augmentation with additive log-normal shadowing; (ii) a RF regression trained per-PCI; (iii) a GPR with a SE kernel; and (iv) a GPR with a RQ kernel; (v) an MLP with three hidden layers; (vi) a CGAN; (vii) a ReFlow model, representing state-of-the-art diffusion-based generation; and (viii) a NF model. All radio feature models are trained on region-specific training sets and evaluated on held-out test sets using the MedAE on per-PCI RSRP values. This evaluation is conducted independently for each geographic region. The reported values in Tab. IV represent the average MedAE over ten independent runs, with associated 95% CI.
The results demonstrate that all models achieve relatively low MedAE values, indicating effective RSRP estimation across diverse urban and semi-urban regions. The KNN model consistently yields the lowest MedAE across all regions, particularly excelling in the airport and city center zones, which contain the highest number of UEs. In contrast, the GPR models show slightly worse performance overall. The RQ variant consistently outperforms the standard GPR across all regions, as it incorporates a more flexible kernel capable of modeling multi-scale signal variations. RF exhibits moderate performance, with accuracy generally falling between KNN and GPR. Moreover, the experimental results reveal a relevant trend: lightweight, instance-based methods (specifically KNN) consistently outperform complex deep generative architectures (such as CGAN, ReFlow, and NF) in RSRP estimation. This performance gap is primarily driven by the context-agnostic nature of the problem, where the input space is restricted to low-dimensional spatial coordinates (latitude, longitude) without auxiliary environmental metadata. While deep learning models typically excel at extracting hierarchical features from high-dimensional structured data, they often struggle to model the high-frequency spatial variations inherent to radio propagation (e.g., local shadowing and multipath) when constrained to such sparse inputs, which is a phenomenon often linked to the spectral bias of neural networks [32]. In contrast, KNN operates as a non-parametric learner with a strong local inductive bias. By avoiding assumptions about global function smoothness, it effectively captures the localized and irregular nature of the radio environment, proving that in the absence of geometric features, local interpolation remains superior to deep functional approximation.
@@ VI-B Results on Radio Features Augmentation
[…]
Table IV: MedAE of RSRP prediction for considered approaches in each reference region.
VI-C Results on Multi-cell Fingerprinting-based Positioning
In this section, we evaluate the end-to-end impact of the proposed mobile data augmentation architecture on multi-cell fingerprinting-based positioning. Starting from the region-specific reference database, we compare positioning performance obtained with the original operator-collected fingerprints against databases augmented with synthetic samples produced by the spatial and radio augmentation stages. The positioning algorithm under test is the wKNN estimator introduced in Sec. V-A; results are reported for all four MDT reference scenarios.
We compare our proposed two-stage architecture KDE-KNN against seven alternative two-stage configurations: KDE-RF, KDE-GPR (SE/RQ), KDE-MLP, KDE-CGAN, KDE-ReFlow, and KDE-NF. For each MDT scenario, we evaluate a set of augmentation rates $A\in\{1,5,10,20,30\}$, where $A$ denotes the multiplicative increase of the fingerprinting database relative to the original ($A=1$ corresponds to no augmentation). For every combination of scenario, augmentation rate and architecture, we perform ten independent runs. In the remanding of this section, we report average MedAE values for positioning with 95% CI.
@@ VI-C Results on Multi-cell Fingerprinting-based Positioning
    City Center Area (Tab. V). In the city center scenario, which features the highest measurement density and the smallest spatial extent, the wKNN estimator applied to the original MDT baseline attains a median positioning error of 20.0820.08\,m. With the proposed KDE-KNN pipeline and an augmentation factor $A=20$, the median error decreases to 17.8017.80\,m, corresponding to an $\approx 11.4\%\$relative improvement. Although the absolute reduction ($\approx 2.3\,$m) is modest, it is noteworthy given the strong baseline performance. Moreover, KDE-KNN consistently outperforms the alternative two-stage configurations, which yield smaller or no gains; in particular, the KDE-RF variant may be detrimental, since RF-based RSRP predictions can produce augmented fingerprints that degrade positioning accuracy.
@@
    Stadium Area (Tab. VI).
    Augm. rate $A$ KDE-KNN KDE-RF KDE-GPR KDE-GPR(RQ) KDE-MLP KDE-CGAN KDE-ReFlow KDE-NF
@@
    Table V: Impact of data augmentation rates on positioning across increasing synthetic data ratios in the city center scenario.
    Augm. rate $A$ KDE-KNN KDE-RF KDE-GPR KDE-GPR(RQ) KDE-MLP KDE-CGAN KDE-ReFlow KDE-NF
@@
    Table VI: Impact of data augmentation rates on positioning across increasing synthetic data ratios in the stadium scenario.
    In the stadium scenario, characterized by a high measurement density but a very confined spatial extent and an intermediate number of PCIs, the proposed KDE-KNN pipeline yields substantial gains. With augmentation factor $A=20$, the median positioning error drops from $36.70$ m to $28.32$ m, an absolute improvement of $8.38$ m and a relative reduction of $\approx 23\%$. By contrast, the alternative two-stage configurations provide only marginal or no improvements across augmentation levels.
  3. 3.
    Airport Area (Tab. VII).
    Augm. rate $A$ KDE-KNN KDE-RF KDE-GPR KDE-GPR(RQ) KDE-MLP KDE-CGAN KDE-ReFlow KDE-NF
@@
    Table VII: Impact of data augmentation rates on positioning across increasing synthetic data ratios in the airport scenario.
@@
    Augm. rate $A$ KDE-KNN KDE-RF KDE-GPR KDE-GPR(RQ) KDE-MLP KDE-CGAN KDE-ReFlow KDE-NF
@@
    Table VIII: Impact of data augmentation rates on positioning across increasing synthetic data ratios in the highway scenario.
    In the highway scenario, where UE measurements are highly concentrated along the linear roadway (i.e., samples cluster on the highway) but sparse elsewhere across the largest geographic extent and subject to high mobility, augmentation delivers the largest benefit. With the proposed KDE-KNN pipeline and augmentation factor $A=30$, the median positioning error falls from $115.51$ m to $79.82$ m, a relative reduction of $\approx 30\%$. This substantial improvement indicates that realistic augmentation can compensate for the anisotropic and sparse spatial sampling typical of highway measurements; by contrast, the alternative two-stage configurations produce markedly smaller gains and are consistently outperformed by KDE-KNN across augmentation levels.
These findings lead to three relevant conclusions: (i) the KDE-KNN pipeline consistently improves positioning accuracy across all evaluated scenarios, with the largest relative gains in spatially sparse or structurally complex regions; (ii) the alternative two-stage baselines fail to match the proposed approach across the tested augmentation rates; (iii) performance gains saturate beyond a region-specific augmentation factor, so adding more synthetic samples does not by itself guarantee better accuracy, and poorly matched synthetic fingerprints can even degrade it; and (iv) in high-density scenarios (e.g., city center or stadium), the saturation in terms of positioning performance occurs at low augmentation rates and converges to low values, as UE density is already high in the training set, as shown in Table II. Conversely, in the highway scenario, characterized by high mobility and extremely low spatial density, the impact of data sparsity is significant. Consequently, this region requires much higher augmentation rates before saturation is reached and overall positioning algorithms perform worse.
VI-D Statistical Model Comparison